OpenStudy (anonymous):

find tan4x? tanx=1/5

hartnn (hartnn):

@TheMightyThumb Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) you can use the formula \(\large \tan2x=\dfrac{2\tan x}{1-\tan^2x}\) to get tan 4x , once to find tan 2x from tan x and once more to find tan 4x from tan 2x

OpenStudy (anonymous):

Wait , so whats the answer

hartnn (hartnn):

you can use the formula to get the answer by yourself! first find tan 2x using that formula, what u get ?

hartnn (hartnn):

Aslo, its against our code of conduct to give out direct answers, you can check it out : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

I got 5/2 for tan2x. What should I do next?

hartnn (hartnn):

tan 2x = 2/5 / (1-1/25) = 2/5 /(24/25) = 2*25/ (24*5) = 5/12 did you mean 5/12 and not 5/2 ??

hartnn (hartnn):

anyways, next will be \(\large \tan4x=\dfrac{2\tan 2x}{1-\tan^22x}\)

OpenStudy (anonymous):

Oh mb. Couldnt read my writing lol

hartnn (hartnn):

using same formula and tan2x = 5/12, find tan 4x :) can you ?

OpenStudy (anonymous):

So just plug in 5/12 for tan2x?

hartnn (hartnn):

yes, absolutely :)

OpenStudy (anonymous):

I got 120/119. Is that right? Thanks

hartnn (hartnn):

That is absolutely correct! good work :) welcome ^_^

OpenStudy (anonymous):

We need more people like you in the world. Thanks ^-^

hartnn (hartnn):

you'll find them in OpenStudy itself :) many more ;) If you need any help browsing this site, you can ask me/message me :) or any other ambassador with an encircled A before their name. Have a good day :)

OpenStudy (anonymous):

I will definitely come here more often. Thanks :)